OpenStudy (anonymous):

17 less than twice a number is equal to the number. what is the number?

OpenStudy (anonymous):

\[2c-17=c\]\[2c-c=17\]\[c=17\]

OpenStudy (anonymous):

17 is the answer 2x-17=x subtract 17 from both sides ot get 2x-x=17 then sutract x from 2x and you get x which leaves you with x=17

OpenStudy (anonymous):

oh ok thakns